OpenStudy (anonymous):

Jess observed that 60% of the people at a mall on a particular day shopped for clothes. If 2500 people at the mall did not shop for clothes that day, the number of people who shopped for clothes that day was ______. (only put numeric values, no other symbols)

OpenStudy (anonymous):

@sleepyjess please help

OpenStudy (anonymous):

@timo86m @dan815 help

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

@carlyakaj

OpenStudy (anonymous):

PLEASE HELP

jigglypuff314 (jigglypuff314):

100% - 60% = 40% so 40% = number of people that did Not shop = 2500 people

OpenStudy (anonymous):

part of the question says (only put numeric values, no other symbols) so would i just put 40 with out the %

jigglypuff314 (jigglypuff314):

I wasn't quite done, sorry... solve for x (stands for the 100%) x * 0.4 = 2500

OpenStudy (anonymous):

are you good at algebra? would you be able to help me on more questions

jigglypuff314 (jigglypuff314):

yeah, I can help with more :)

OpenStudy (anonymous):

ok

jigglypuff314 (jigglypuff314):

once you get x from the x * 0.4 = 2500 do x * 0.6 = (answer you want)

OpenStudy (anonymous):

so the answer would be 0.6

jigglypuff314 (jigglypuff314):

not quite :) what do you get when you solve for x for 0.4x = 2500 ?

OpenStudy (anonymous):

you would get either 6250 or 0.00016 i think

jigglypuff314 (jigglypuff314):

x = 6250 :) then 6250 times 0.6 = (answer you want)

OpenStudy (anonymous):

so the answer would be 3750

jigglypuff314 (jigglypuff314):

correct! :)

OpenStudy (anonymous):

thanks is that the final answer

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Beluga Fun Tours: Age Ticket Price Under 3 years free 3 to 12 years $34 Over 12 years $45 On a certain day, the company took 103 people on whale watching trips. There were 64 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under 3 years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (64 - x)34 + (103 - 64) ⋅ 45 = C (103 + x)34 + (103 -(64 + x)) ⋅ 45 = C (103 - x)34 + (103 - 64) ⋅ 45 = C (64 + x)34 + (103 -(64 + x)) ⋅ 45 = C

OpenStudy (anonymous):

what about thhis one

jigglypuff314 (jigglypuff314):

gimme a minute :)

OpenStudy (anonymous):

ok

jigglypuff314 (jigglypuff314):

ok... so the number of "over 12 years" = 103-64 makes sense? and the cost for "over 12 years" = 45 so part of the equation has to include (103-64)45

OpenStudy (anonymous):

ok i undrestand then what

jigglypuff314 (jigglypuff314):

then 64 - x = number of "3 to 12 years" and the cost of "3 to 12 years" = 34 so it would also include the part (64 - x)34

OpenStudy (anonymous):

so then you would do (103-61) * 45=c

jigglypuff314 (jigglypuff314):

not quite :) take a look at the given choices and combine the two parts that I gave :)

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

would it be the 1st answer (64 - x)34 + (103 - 64) ⋅ 45 = C

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

The width of a rectangle is 12 units. Can the perimeter x of the rectangle can be 60 units when its length y is 18 units? No, the rectangle cannot have x = 60 and y = 18 because x + y is less than 60. No, the rectangle cannot have x = 60 and y = 18 because x + y is less than 24. Yes, the rectangle can have x = 60 and y = 18 because x = 12 + 2y. Yes, the rectangle can have x = 60 and y = 18 because x = 24 + 2y.

OpenStudy (anonymous):

im on this practice test and i just want help to be ready for the pretest

jigglypuff314 (jigglypuff314):

x = 2(12) + 2(y) if y=18 plug it in and see if x=60

OpenStudy (anonymous):

yes it equals 60

jigglypuff314 (jigglypuff314):

then "Yes, the rectangle can have x = 60 and y = 18 because x = 24 + 2y"

OpenStudy (anonymous):

Equation 1 |4x – 3|− 5 = 4 Equation 2 |2x + 3| + 8 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = -4, 1. The solutions to equation 1 are x = 3, -1.5 and equation 2 has no solution. The solutions to equation 1 are x = 3, -1.5 and equation 2 has solutions x = -4, 1.

jigglypuff314 (jigglypuff314):

well, the way would would solve an absolute value equation would be |4x – 3|− 5 = 4 isolate abs value by add 5 to both sides +5 +5 |4x - 3| = 9 which can be separated into 4x - 3 = 9 and 4x - 3 = -9 then solve for x for both

OpenStudy (anonymous):

x=-1.5

jigglypuff314 (jigglypuff314):

absolute value equations separate into two equations so that x can equal two values

OpenStudy (anonymous):

so would it be " The solutions to equation 1 are x = 3, -1.5 and equation 2 has no solution."

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4 Which of the following is a correct next step to solve the equation? x + 2 = -4 -x – 2 = 4 x + 2 = 4 x - 2 = -4

jigglypuff314 (jigglypuff314):

|x - 2| = 4 can be separated into x - 2 = 4 and x - 2 = -4

OpenStudy (anonymous):

opps sorry it would be x-2=-4

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 1) greater than or equal to -12? x greater than or equal to -5 x less than or greater to -5 -5 less than or greater to x A number line with a closed circle on negative 5 and shading to the right

jigglypuff314 (jigglypuff314):

solve for x as would would for an equation. 2(x-1) > -12 distribute the 2 add 2 to both sides divide both sides by 2 and you get...

OpenStudy (anonymous):

x greater than or equal to -5

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

Choose the missing step in the given solution to the inequality -x - 10 > 14 + 2x. -x - 10 > 14 + 2x -3x -10 > 14 x < -8 -3x > 24 -3x < 4 -3x > -4 8 < x

jigglypuff314 (jigglypuff314):

that weird >,< there's multiple typos?

jigglypuff314 (jigglypuff314):

oh wait... or are the last four options?

OpenStudy (anonymous):

the last 4 options

jigglypuff314 (jigglypuff314):

from -3x -10 > 14 you'll need to add 10 to both sides so once you add 10 to both sides, what do you get?

OpenStudy (anonymous):

-3x>24

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 3(2x - 1) greater than or equal to 4(2x - 3) - 3? x less than or greater to 6 6 greater than or equal to x A number line with a closed circle on 6 and shading to the left A number line with a closed circle on 6 and shading to the right

jigglypuff314 (jigglypuff314):

sorry, I just realized this question said "Not" Which of the following is not a way to represent the solution of the inequality 2(x - 1) greater than or equal to -12? x greater than or equal to -5 x less than or greater to -5 <- so it would be this one -5 less than or greater to x A number line with a closed circle on negative 5 and shading to the right

OpenStudy (anonymous):

OK

jigglypuff314 (jigglypuff314):

to solve this inequality 3(2x - 1) > 4(2x - 3) - 3 distribute on both sides add like terms subtract 8x from both sides add 3 to both sides divide both sides by -2 and don't forget to flip the sign :)

OpenStudy (anonymous):

x<6

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

so its x less than or greater to 6

jigglypuff314 (jigglypuff314):

umm it's asking "Which of the following is NOT a way"

OpenStudy (anonymous):

oh sorry it kicked me out and so it would be 6 greater than or equal to x

jigglypuff314 (jigglypuff314):

x<6 is same as 6>x

OpenStudy (anonymous):

A number line with a closed circle on 6 and shading to the right

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

Abbey wants to use her savings of $1325 to learn yoga. The total charges to learn yoga include a fixed registration fee of $35 and a monthly fee of $50. What is the maximum number of months for which Abbey can learn yoga with her savings? 24 25 26 27

jigglypuff314 (jigglypuff314):

solve the inequality 50x + 35 < 1325

OpenStudy (anonymous):

26

jigglypuff314 (jigglypuff314):

you would end up with x < 25.8 so round down :)

OpenStudy (anonymous):

so it is 26 or 24

jigglypuff314 (jigglypuff314):

what is the integer less than 25.8?

OpenStudy (anonymous):

27

jigglypuff314 (jigglypuff314):

less than 25.8?

jigglypuff314 (jigglypuff314):

would be 25

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 1 less than or greater to 12 or 2x greater than or equal to 30. A number line with an open circle on 13, shading to the left, and an open circle on 15, shading to the right. A number line with a closed circle on 13, shading to the left, and a closed circle on 15, shading to the right. A number line with an open circle on 13, an open circle on 15, and shading in between. A number line with a closed circle on 13, a closed circle on 15, and shading in between.

jigglypuff314 (jigglypuff314):

solve for x for x - 1 < 12

OpenStudy (anonymous):

x<13

jigglypuff314 (jigglypuff314):

then solve for x for 2x > 30

OpenStudy (anonymous):

x>15

jigglypuff314 (jigglypuff314):

then they are both closed hole cuz of the "or equal to"

OpenStudy (anonymous):

so it would be A number line with a closed circle on 13, shading to the left, and a closed circle on 15, shading to the right.

OpenStudy (anonymous):

3x + 1 > 10 and 4x less than or greater to 20. A number line with an open circle on 3, a closed circle on 5, and shading in between. A number line with an open circle on 3, shading to the left, and a closed circle on 5, shading to the right. A number line with a closed circle on 3, shading to the left, and an open circle on 5, and shading to the right. A number line with a closed circle on 3, an open circle on 5, and shading in between.

jigglypuff314 (jigglypuff314):

solve for both separately 3x + 1 > 10 4x < 20

jigglypuff314 (jigglypuff314):

I've gtg x > 3 x <_ 5 open hole at 3 because has no "or equal to" closed hole at 5 because has "or equal to" shaded in between